OpenStudy (math_bender_15):

The container that holds the water for the football team is 1/3 full. After pouring in 6 gallons of water, it is 1/2 full. How many gallons can the container hold?

OpenStudy (studyhard):

Let amount of water container can hold be x. 1/3x+6=1/2x. 1/2x-1/3x=6. 1/6x=6. x=36. Therefore, the container can hold 36 gallons of water.